A valószínűségszámítás és matematikai statisztika gyakorlaton való részvétel

dátum:  IX. 9. IX. 16. IX. 23. IX. 30. X. 7. X. 14. XI. 4.

diákok:

Agócs László Péter  +  +  −  +  +  +  +
Arany Gábor  +  +  +  +  +  +  −
Bálint Levente  +  +  +  +  +  −  +
Csipkés László  +  +  −  +  +  +  +
Eszenyi László  −  +  −  +  +  +  +
Ferenczik Zoltán  +  +  +  +  +  +  +
Fogarasi Edina  +  −  +  +  +  +  +
Gerják Balázs  +  +  +  +  +  +  +
Juhász Tibor  +  +  +  +  +  +  +
Kis Bertalan  +  +  +  +  +  +  +
Kocsis Gergely  +  +  +  +  +  +  +
Kovács György  +  +  −  +  +  +  +
Lázár Gábor  +  −  +  +  +  +  +
Nagy Csaba  +  +  −  +  +  +  +
Paluska Mihály  +  +  +  +  +  +  +
Pári Viktor  +  +  +  +  −  +  +
Pelejtei Gergely  +  −  +  +  +  +  +
Sipos Nándor  +  +  +  +  +  +  +
Szakácsi János  +  +  +  +  +  +  +
Szalontai Gábor  −  +  +  +  +  +  +
Szalontai József  +  +  +  +  +  +  +
Tóth János  +  +  +  +  +  +  +
Tóth Tamás  +  +  +  +  +  +  +
Virág Tamás  +  +  +  +  +  +  +
Zabos Katalin  +  +  +  +  −  +  −

dátum:  IX. 9. IX. 16. IX. 23. IX. 30. X. 7. X. 14. XI. 4.

diákok:

Dancsi József  +  −  +  +  +  +
Emődi Szabolcs Árpád  +  +  +  +  +  +
Fülöp Arnold  +  +  −  +  +  +
Hamza Péter  +  +  −  −  +  +
Hank Sándor  +  +  −  −  +  +
Lörinczi József  +  +  +  +  +  +
Molnár Angéla  +  +  +  +  +  +
Molnár Tibor  −  +  +  −  +  +
Nagy Gábor  +  −  +  +  +  +
Nagy Zsolt  +  +  +  +  +  +
Páli Gábor János  +  +  +  +  +  +
Paper Andrea  +  −  +  −  +  +
Papp Gyula  +  +  +  +  +  +
Szabó Zsolt  +  +  +  +  +  +
Szevelényi Attila  +  +  +  +  +  +
Tilki Csaba  −  +  −  +  +  +
Tordai Jácint  +  +  +  −  +  +

dátum:  XI. 11. XI. 18. XI. 25. XII. 2. XII. 9. XII. 16.

diákok:

Agócs László Péter  +
Arany Gábor  +
Bálint Levente  +
Csipkés László  +
Eszenyi László  +
Ferenczik Zoltán  +
Fogarasi Edina  +
Gerják Balázs  +
Juhász Tibor  +
Kis Bertalan  +
Kocsis Gergely  +
Kovács György  +
Lázár Gábor  −
Nagy Csaba  +
Paluska Mihály  +
Pári Viktor  +
Pelejtei Gergely  −
Sipos Nándor  +
Szakácsi János  +
Szalontai Gábor  +
Szalontai József  −
Tóth János  +
Tóth Tamás  +
Virág Tamás  +
Zabos Katalin  +

dátum:  XI. 11. XI. 18. XI. 25. XII. 2. XII. 9. XII. 16.

diákok:

Dancsi József  −
Emődi Szabolcs Árpád  +
Fülöp Arnold  +
Hamza Péter  +
Hank Sándor  −
Lörinczi József  +
Molnár Angéla  +
Molnár Tibor  +
Nagy Gábor  +
Nagy Zsolt  +
Páli Gábor János  +
Paper Andrea  −
Papp Gyula  +
Szabó Zsolt  +
Szevelényi Attila  +
Tilki Csaba  +
Tordai Jácint  −